Please help me with this question:
What is the equation of a line, g(x), that is perpendicular to f(x)=1/2x+2 and passes through (0,2)?
a) g(x)=-4x+2
b) g(x)=-4x-2
c) g(x)=-4x-8
d) g(x) can not be determined with this information

Answer by Alan3354(69443)

The slope of f(x) is 1/2, as it's already in the for y=mx+b
The slope of g(x) is -1/2, as it's perpendicular.
(y - Y) = m(x - X) (X,Y) is point (0,2)
y - 2 = -1/2(x - 0)
y - 2 = -x/2
2y - 4 = -x
x+2y = 4